# Collective order boundedness of sets of operators between ordered vector spaces

### v2

@@ Abstract
It is proved that: each collectively order continuous set of operators between Archimedean ordered vector spaces with generating cones is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
@@ 1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 7, 8, 9]. The present paper generalizes results of [1, 7, 15] to the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator between Archimedean ordered vector spaces is order bounded; 2) each order to norm continuous operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL). Throughout the text, vector spaces are real, operators are linear, symbol ${\cal L}(X,Y)$ (${\cal L}_{+}(X,Y)$, ${\cal L}_{ob}(X,Y)$, ${\cal L}_{oc}(X,Y)$, ${\cal L}_{onb}(X,Y)$) stands for the space of all (positive, order bounded, order continuous, order to norm bounded) operators from $X$ to $Y$. A symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
@@
We shall use the following OVS version of [7, Definition 2.1]. Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
@@
The following definition extends [7, Definition 2.1] to OVSs.
@@
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an OVS to an Archimedean OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded. Section 3 is devoted to applications of obtained results to commutative operator semigroups on OVSs.
@@ 2 Main results
We are going to prove the following results which generalize the results obtained recently in [7, 15] in the vector lattice setting to the OVS setting.
@@
    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS with a generating cone is collectively order bounded.
@@
    Collectively r-continuous sets coincides with collectively order bounded sets when the OVS in the range has a generating cone.
@@
We begin with generalization of the first part of [7, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
@@
Let $X$ and $Y$ be OVSs with generating cones and suppose $X$ is Archimedean. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
@@
Let ${\cal T}\in\text{\bf L}_{oc}(X,Y)$. It is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$. Take $[0,b]\subseteq X$. Following [1, Theorem 2.1], consider the lexicographically directed set =×${\cal I}=\mathbb{N}\times[0,b]$: $(m,z)\geq(n,z^{\prime})$ if $m>n$ or $m=n$ and $z\geq z^{\prime}$, and let $x_{(k,z)}=\frac{1}{k}z$. Since $X$ is Archimedean, $\frac{1}{n}b\downarrow 0$. It follows from $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$ for $(m,z)\geq(n,b)$ that $x_{(m,z)}\operatorname{\xrightarrow[]{o}}0$.
@@
Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $Y$ with generating cone is order bounded.
The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Corollary 2.1 even in the VL setting.
@@
Next, we generalize [15, Theorem 5.1] and [7, Theorem 2.2] to OVSs.
@@
Let $X$ and $Y$ be OVSs with generating cones then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
@@
Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. In order to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $[0,b]\subseteq X$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{r}}0$ in $X$.
@@
Now, suppose $Y=Y_{+}-Y_{+}$. Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$. Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$, there exist $w_{1},w_{2}\in Y_{+}$ satisfying $Tu\in[-w_{1},w_{1}]$ and $T(u+nx_{\alpha})\in[-w_{2},w_{2}]$ for all $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. Then $nTx_{\alpha}\in[-(w_{1}+w_{2}),(w_{1}+w_{2})]$ (and hence ±$\pm Tx_{\alpha}\leq\frac{w_{1}+w_{2}}{n}$) for $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. ∎
@@
The following result generalizes [7, Theorem 2.1] to OBSs.
@@ 3 Some applications to operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [13, p.75]).
@@
Finally, we include the following convergence theorem.
@@
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [13, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
@@
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs.
The next two lemmas follow directly from [7, Theorems 2.1 and 2.2] and from Theorems 2.1, 2.2, and 2.3.
@@
Every collectively order to norm bounded set of operators on an (OBS) with a closed generating cone is norm bounded.
@@ 4. $iv)$
In recent works [10, 11, 12] relatively uniformly continuous semigroups on VLs were deeply studied. The use of collectively qualified semigroups is to rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.

### v3

@@ Abstract
It is proved that: each collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
@@ 1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 8, 9, 10]. The present paper generalizes previous results of [1, 6, 8, 16] to collectively qualified set of operators in the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm continuous operator on an ordered Banach space with a closed generating cone is bounded.
We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL). Throughout the text, vector spaces are real, operators are linear, a symbol ${\cal L}(X,Y)$ stands for the space of operators from $X$ to $Y$, and a symbol $x_{\alpha}\downarrow 0$ for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
@@
We shall use the following OVS version of [8, Definition 2.1]. Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
@@
The following definition extends [8, Definition 2.1] to OVSs.
@@
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded. Section 3 is devoted to applications of obtained results to operator semigroups on OVSs.
@@ 2 Main results
We are going to prove the following results which generalize the results obtained in [6, 8, 16] in the vector lattice setting to the OVS setting.
@@
    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded.
@@
    Collectively r-continuous sets of operators between OVS with generating cones agree with collectively order bounded sets.
@@
We begin with generalization of the first part of [8, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
@@
If $X$ is an Archimedean OVS with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for each OVS $Y$.
@@
Let ${\cal T}\in\text{\bf L}_{oc}(X,Y)$. Since $X_{+}-X_{+}=X$, it is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$. Take $[0,b]\subseteq X$. Following [1, Theorem 2.1], consider the lexicographically directed set =×${\cal I}=\mathbb{N}\times[0,b]$: $(m,z)\geq(n,z^{\prime})$ if $m>n$ or $m=n$ and $z\geq z^{\prime}$, and let $x_{(k,z)}=\frac{1}{k}z$. Since $X$ is Archimedean, $\frac{1}{n}b\downarrow 0$. It follows from $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$ for $(m,z)\geq(n,b)$ that $x_{(m,z)}\operatorname{\xrightarrow[]{o}}0$.
@@
Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $T$ is order bounded.
The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and Corollary 2.1 even in the VL setting.
@@
Next, we generalize [6, Theorem 4.6.2], [16, Theorem 5.1], and [8, Theorem 2.2] to OVSs as follows.
@@
Let $X$ and $Y$ be OVSs.
  1. $i)$
    If $X$ has a generating cone then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
  2. $ii)$
    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{rc}(X,Y)$.
The idea of the proof follows the proof of [6, Theorem 4.6.2] with an additional use of the special net $x_{(k,z)}=\frac{1}{k}z$ form [1, Theorem 2.1] like in the proof of Theorem 2.1.
@@ 2. $ii)$
$i)$   Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. Since $X_{+}-X_{+}=X$, for ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $b\in X_{+}$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{r}}0$ in $X$.
@@
$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$. Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$, there exist $w_{1},w_{2}\in Y_{+}$ satisfying $Tu\in[-w_{1},w_{1}]$ and $T(u+nx_{\alpha})\in[-w_{2},w_{2}]$ for all $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. Then $nTx_{\alpha}\in[-(w_{1}+w_{2}),(w_{1}+w_{2})]$ (and hence ±$\pm Tx_{\alpha}\leq\frac{w_{1}+w_{2}}{n}$) for $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. ∎
@@
The following result generalizes [8, Theorem 2.1] to OBSs.
@@ 3 Some applications to operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [14, p.75]).
@@
We include the following convergence theorem.
@@
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [14, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
@@
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs. The next two lemmas follow directly from Theorems 2.1, 2.2, and 2.3.
@@
Every collectively order to norm bounded set of operators on an OBS with a closed generating cone is norm bounded.
@@ 4. $iv)$
In recent works [11, 12, 13] relatively uniformly continuous semigroups on VLs were deeply studied. The use of collectively qualified semigroups is too rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.